OpenStudy (laddiusmaximus):

Im a little fuzzy on the setup on this question: Find the center and radius of the given circle. x^2 + y^2 = 16

hartnn (hartnn):

the center and radius of the given circle, \( x^2 + y^2 = r^2 \) is C=(0,0) , radius =r. can you find r ?

OpenStudy (laddiusmaximus):

that would be 4^2 right? love your profile pic btw.

hartnn (hartnn):

right, so r=radius =4 , center was (0,0) and thanks :)

OpenStudy (laddiusmaximus):

so is that it? that's all I need to do?

hartnn (hartnn):

SORRY FOR LATE REPLY AND YES.